OpenStudy (anonymous):

(x+8)^2+(y-4)^2=81 whats the center and the radius? im confused on how to do this any help I will fan and medal

OpenStudy (shamim):

actually the equation of a circle is\[x ^{2}+y ^{2}=a ^{2}\]

OpenStudy (shamim):

u hv to compare ur given equation with my equation

OpenStudy (shamim):

obviously \[a ^{2}=81\]

OpenStudy (shamim):

\[a=9\]

OpenStudy (shamim):

a=radius of a circle

OpenStudy (anonymous):

the center and radius eq of the circle is \[ (x – h)^2 + (y – k)^2 = r^2\] with the center being at point (h,k) and the radius being r

OpenStudy (shamim):

and in my equation the center will b (0,0)

OpenStudy (shamim):

so the center will b\[x+8=0\]

OpenStudy (shamim):

\[x=-8\]

OpenStudy (anonymous):

so what would be the center and radius in your problem @kittykat96?

OpenStudy (shamim):

and\[y-4=0\]

OpenStudy (shamim):

\[y=4\]

OpenStudy (anonymous):

actually the equation of a circle with center \((h,k)\) and radius \(r\) is \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (shamim):

so center is (-8,4)

OpenStudy (anonymous):

yup i wrote that but may have worded it wrong @satellite73

OpenStudy (anonymous):

ok so the center is (-8,4) and the radius is 9? im really confused right now so many people trying to answer at once lol but I think I got it

OpenStudy (anonymous):

yup that is the answer but use the formula @satellite73 and i put up when you want to solve for the center and radius r

OpenStudy (anonymous):

yes, you have it

OpenStudy (anonymous):

ok thank you guys